What is the energy in kJ of a mole of photons with a wavelength of 440 nm?
Answer #1

Given:

lambda = 440 nm = 4.4*10^-7 m

Find energy of 1 photon first

use:

E = h*c/lambda

=(6.626*10^-34 J.s)*(3.0*10^8 m/s)/(4.4*10^-7 m)

= 4.518*10^-19 J

This is energy of 1 photon

Energy of 1 mol = energy of 1 photon * Avogadro's number

= 4.518*10^-19*6.022*10^23 J/mol

= 2.721*10^5 J

= 2.721*10^2 KJ

Answer: 272 KJ